m=(m=( -7+(-2 )/2, 3+(-3)/2


1
Expert's answer
2022-01-24T17:12:15-0500

"m=(-7+\frac{(-2)}{2}, 3+\frac{(-3)}{2})=(-8, \frac{3}{2})"